OpenStudy (mikee):

The following table lists prices for various packages of Tide laundry detergent powder. Find the unit price of each package. Then determine which one is the best buy. 33oz cost $3.69 87oz cost $6.22 131oz cost $10.99 263oz cost $17.99

OpenStudy (anonymous):

263 oz. cost $17.99

OpenStudy (dumbcow):

unit price means cost per oz 3.69/33 = .11 6.22/87 = .07 10.99/131 = .08 17.99/263 = .068 so you see the lowest and best buy is 263 oz for 17.99

OpenStudy (mikee):

Ty Mowrey and Dumbcow